- Find the volume of the soud that lies within both the cyunder x2+y2-1 and the sphere x++y24724
Answer #1

We have x² + y² =)82 2 2 and a²+1²+2²= 4 and p²4 2 ² - 4 3) 0 2² 4-02 -) Z=14-12 trow, m warries from 1 we have. Brot the fou2TT ป (-233) do -243) 40 3 D Lๆ 16 3 243) Te]: (270)( - 2J3) ว่า IC - 41(N3 3 32 ... Required volume in 32 T - 4ง3